# Level-Set Parameters: Novel Representation for 3D Shape Analysis

### v1

@@ 6 Conclusion
Limitations. Level-set parameters summarize the overall topology of 3D shapes and can be applied to global feature learning of continuous shapes. However, they are not suitable for learning local features of 3D shapes, as there is no correspondence between the local structures of the shape and subsets of the level-set parameters. Researchers have explored local modulation vectors [76] to improve the image classification performance of continuous representations, but with limited success. Mixtures of neural implicit functions [77] may offer potential for enhanced local encoding in the continuous representations. In addition, distinguishing the level-set parameters into pose-dependent and independent subsets, and making them learnable from arbitrarily posed point clouds, could be important for continuous shape representations and analysis in the future.
@@ A.1 Tensors of Level-set Parameters
We show the 8-layer SDF network with skipping concatenation in Fig. 6. The resulting level-set parameters $\Theta_{1},\Theta_{2},\Theta_{3}$ have dimensions $256{\times}4$, $6{\times}256{\times}257$, and $1{\times}257$, respectively. We use the proposed hypernetwork $h_{\boldsymbol{\phi}}$ to generate the first layer parameters for surface transformation.
@@ B.1 Geometric SDF Initialization with the Hypernetwork
The normalized variable $\hat{Z}{=}\frac{Z}{\sqrt{\sum_{i}{\beta_{i}^{2}}}}\sim\mathcal{N}(0,1)$ [78]. We normalize each element in $\mathbf{z}$ as $\hat{z}_{j}{=}\frac{z_{j}}{\sqrt{\sum_{i}{B_{ij}^{2}}}}$ to obtain a normalized vector $\hat{\mathbf{z}}$ where $\hat{z}_{j}\sim\mathcal{N}(0,1)$.
Given the normalized vector $\hat{\mathbf{z}}$ which strictly follows a standard normal distribution, we design two different branches, FC1 and FC2, in the last layer of $h_{\boldsymbol{\phi}}$ to initialize the weight and bias parameters in $\{\theta_{1}^{mn}\}$ according to SAL [58] . We note that the approximate Gaussian Process in [79] does not help in generating outputs that follow the standard normal distribution for the required initializations.
@@
The proposed hypernetwork emphasizes importance of the first SDF layer while allowing the SDF parameters of the other layers to be shared across different poses, substantially reducing the neural parameter size of $h_{\boldsymbol{\phi}}$. The introduction of the latent matrices $\{\mathbf{Y}^{mn}\}$ is important as it enables the hypernetwork to satisfy the geometric initializations of SDF network. Note that parameters of the unconditioned layers (layer 2-8) in SDF are initialized the same as in SAL [58].
@@ Appendix C Implementation Details
Hypernetwork $h_{\boldsymbol{\phi}}$. In the first stage of our dataset construction, we train HyperSE3-SDF with a batch size of 50 for 50000 epochs. We utilize the Adam Optimizer [80] with an initial learning rate of 0.001, which exponentially decays at a rate of $\gamma{=}0.998$ every 30 epochs to train the model. This training process takes 3 days on a GeForce RTX 4090 GPU.
@@ Appendix D Pose-Dependent Parameters in $\boldsymbol{\theta}$
We evaluate surface reconstruction using L1 Chamfer Distance (CD1×$\times 100$) and Normal Consistency (NC). To assess overall surface quality across various transformations, we randomly sampled 50 poses, calculating the mean and deviation of these metrics to report performance. Table 5 summarizes the results, showing that the proposed hypernetwork significantly outperforms Baseline and Baseline++. For reference, we provide the corresponding shape quality in our two-stage constructed dataset. The training time for a shape in the second stage of dataset construction is 4 minutes, which is significantly less than the 1 hour required for a hypernetwork without learned initializations.
@@ Appendix E Pose Estimation
Figure 9: Registration comparison between Teaser++ and the proposed method. We report RRE/RTE results. Note that RTE is scaled by ×$\times 100$.
@@
This work focuses on demonstrating the viability of level-set parameters as an independent data modality with transformations enabled in SE(3) for shape representations. It is different from some existing work [81] which exploits neural fields as a self-supervision technique to learn semantic features disentangled from the orientation of explicit representations such as images.

### v2

@@ 6 Conclusion
Limitations. Level-set parameters summarize the overall topology of 3D shapes and can be applied to global feature learning of continuous shapes. However, they are not suitable for learning local features of 3D shapes, as there is no correspondence between the local structures of the shape and subsets of the level-set parameters. Researchers have explored local modulation vectors Bauer et al. (2023) to improve the image classification performance of continuous representations, but with limited success. Mixtures of neural implicit functions You et al. (2024) may offer potential for enhanced local encoding in the continuous representations. In addition, distinguishing the level-set parameters into pose-dependent and independent subsets, and making them learnable from arbitrarily posed point clouds, could be important for continuous shape representations and analysis in the future.
@@ A.1 Tensors of Level-set Parameters
[…]
@@ B.1 Geometric SDF Initialization with the Hypernetwork
The normalized variable $\hat{Z}{=}\frac{Z}{\sqrt{\sum_{i}{\beta_{i}^{2}}}}\sim\mathcal{N}(0,1)$ Murphy (2012). We normalize each element in $\mathbf{z}$ as $\hat{z}_{j}{=}\frac{z_{j}}{\sqrt{\sum_{i}{B_{ij}^{2}}}}$ to obtain a normalized vector $\hat{\mathbf{z}}$ where $\hat{z}_{j}\sim\mathcal{N}(0,1)$.
Given the normalized vector $\hat{\mathbf{z}}$ which strictly follows a standard normal distribution, we design two different branches, FC1 and FC2, in the last layer of $h_{\boldsymbol{\phi}}$ to initialize the weight and bias parameters in $\{\theta_{1}^{mn}\}$ according to SAL Atzmon and Lipman (2020) . We note that the approximate Gaussian Process in Lee et al. (2017) does not help in generating outputs that follow the standard normal distribution for the required initializations.
@@
The proposed hypernetwork emphasizes importance of the first SDF layer while allowing the SDF parameters of the other layers to be shared across different poses, substantially reducing the neural parameter size of $h_{\boldsymbol{\phi}}$. The introduction of the latent matrices $\{\mathbf{Y}^{mn}\}$ is important as it enables the hypernetwork to satisfy the geometric initializations of SDF network. Note that parameters of the unconditioned layers (layer 2-8) in SDF are initialized the same as in SAL Atzmon and Lipman (2020).
@@ Appendix C Implementation Details
Hypernetwork $h_{\boldsymbol{\phi}}$. In the first stage of our dataset construction, we train HyperSE3-SDF with a batch size of 50 for 50000 epochs. We utilize the Adam Optimizer Kingma and Ba (2014) with an initial learning rate of 0.001, which exponentially decays at a rate of $\gamma{=}0.998$ every 30 epochs to train the model. This training process takes 3 days on a GeForce RTX 4090 GPU.
@@ Appendix D Pose-Dependent Parameters in $\boldsymbol{\theta}$
We evaluate surface reconstruction using L1 Chamfer Distance (CD1×100$\times$ 100× 100) and Normal Consistency (NC). To assess overall surface quality across various transformations, we randomly sampled 50 poses, calculating the mean and deviation of these metrics to report performance. Table 5 summarizes the results, showing that the proposed hypernetwork significantly outperforms Baseline and Baseline++. For reference, we provide the corresponding shape quality in our two-stage constructed dataset. The training time for a shape in the second stage of dataset construction is 4 minutes, which is significantly less than the 1 hour required for a hypernetwork without learned initializations.
@@ Appendix E Pose Estimation
Figure 9: Registration comparison between Teaser++ and the proposed method. We report RRE/RTE results. Note that RTE is scaled by ×100$\times$ 100× 100.
@@
This work focuses on demonstrating the viability of level-set parameters as an independent data modality with transformations enabled in SE(3) for shape representations. It is different from some existing work Kwon et al. (2023) which exploits neural fields as a self-supervision technique to learn semantic features disentangled from the orientation of explicit representations such as images.